Left Direct Gregorian Technical Observing with IRTC: UT 2010 March 03

Observer: DMiller JHill (Tucson)
Telescope Operator: WWack
Instrument Support:
Telescope Support:
SW support: MDeLaPena, NCushing (Tucson)

Summary

Telescope was not movable till 5:30 UT because of engineering work.

Open the rest of the night with an estimated 0.8" seeing and clear sky. Also note that we had 1 arcsec images on both IRTC and LBC-Red without doing any active optics.

Performed part of the AO2 testing and all of the BP1 testing using IRTC @ LDG and LBC-Red @ prime. Found a problem with LBC's TCS libraries for TCSSendWavefront. IT #2550 was solved later in the day.

Could only close the guide and Active Optics loop for a few minutes at a time till the AzCam in AGW4 failed. Found that the azcam failure was repeatable and sent an email to Torsten and Mike L and added it to IT #2543. This was later diagnosed to be a failed SSD disk in the PEPSI1 computer.

If the AzCam problem is not resolved today, we cannot perform left direct Gregorgian commissioning tonight (4UT)

Details

All times in UT

Open

05:30 Open to mostly clear sky.

05:41 pointto M5_1246 - hung waiting for PCS. Fortunately Michele is right here.

05:45 Stop and restart PCS and IIF.

05:47 pointto M5_1246 - Rotator seems not to be getting polynomials

05:55 Stop PCS. Click Right/Left. Start PCS.

05:56 pointto M5_1246 ACQUIRE - no star, ptspiral, still no star

06:05 Go off to AZ=120 EL=90 reset encoders, ptreset.

06:09 pointto M5_1246 ACQUIRE gs=0 PAM- - no star

Switch to motor encoders, no help. Back to strip.

06:19 Back to AZ=120

06:21 pointto M5_1311 TelMode="ACQUIRE" gs=0 PAM- no star

06:24 home AGW4 again

No star on IRTC while the probe was parking.

06:28 pointto M5_1311 TelMode="ACQUIRE" gs=0 PAM- ptspiral

Mar 3 06:37:59 tcs4 LBT_left_GCS: AzCam: calculating timeout for image from PEPSI2 with size of 567360 bytes and 2000 ms exposureTime to be 6448 ms

GCS was talking to PEPSI2

06:40 Authorize LUCIFER and then back to IRTC@LDG.

The good news. We are now talking to PEPSI1 in AGW4. The bad news: Now the problem with guide images not arriving from AGW4 has returned.

06:46 stopAGW (4)/ start AGW

06:50 pointto M5_1311 TelMode="ACQUIRE" gs=0 PAM-

Couldn't set full frame ROI. Continuing in the hope the image is large enough anyway.

06:52 pointto M5_1311 TelMode="ACQUIRE" gs=0 PAM- - AGw command status: timed out waiting for probe to reach destination
Resmsg[3]  = error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): -1.14302e-08, 0.203414, 33.3876 mm.
Resmsg[4]  = guidestar list successfully changed but probe couldn't be moved to position.

06:54 Home AGW4 again (taking a long time) We leave Wayne fiddling with that and go off to do something else.

We also note that there is zero servo error being reported on MCSPU pos page. A problem? Or impressive servo tuning?

07:05 AGW4 is homed.

LBC Tests with AO2

06:59 Authorize as LBC

07:02 play OB wt000350 LBC turned itself off. Or maybe the server program crashed???

07:06 Turn on LBC-Red again (Blue is not mounted)

07:08 play OB wt000350 070857

Not everything is broken. We see the LBC-Red pointing star just on the edge of IRTC on left direct! This is good, because it means open-loop pointing is within ~20 arcsec (albeit using the LBC-Blue pointing model for Rigid M2). This bodes well for future co-pointing if Left Prime and Left Direct Gregorian have pointing so close.

Adjusting IE/CA moved the mount as expected. Having IRTC running makes this test very fast.

07:16 Press Stop on LBCGUI to get it out of a funny state where it didn't end the OB. Was this because Blue is off?

07:17 play OB wt000350 x200, BUT OB progress still seems to be counting into negative numbers every couple seconds.

07:20 play OB wt000350 x20, BUT OB progress still seems to be counting into negative numbers every couple seconds. Then LBC server program crashes again.

07:25 play OB wt000350 x5 072509

Crashed again after I pressed STOP.

07:30 play OB wt000350 x100 - 073015 guiding failed
2010/03/03 07:31:15.661000 N R TRACKERS IA       "track_next.exe" program on "lbcrtec.20100303.073108_1.fits" and "lbcrtec.20100303.073108_2.fits" images ...
2010/03/03 07:31:16.240000 W R TRACKERS IA       "track_next.exe" program on "lbcrtec.20100303.073108_1.fits" and "lbcrtec.20100303.073108_2.fits" images failed with retcode:18 >no good source found<
2010/03/03 07:31:16.286000 N R TRACKERS IA       doing track NEXT processing completed
2010/03/03 07:31:16.300466 W R TRACKERS          IA System error [src/trackers/trackers.c:1654]
2010/03/03 07:31:16.646000 S R TRACKERS CCDCTRL  electronics temperature: 36.1 [Celsius]

This time it disconnected from TCS rather than crashing.

2010/03/03 07:34:21.036194 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan303" skipped
2010/03/03 07:34:21.036269 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan304" skipped
2010/03/03 07:34:21.036344 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan305" skipped
2010/03/03 07:34:21.036419 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan306" skipped
2010/03/03 07:34:21.036494 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan307" skipped
2010/03/03 07:34:21.036568 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan308" skipped
2010/03/03 07:34:21.036643 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan309" skipped
2010/03/03 07:34:21.036718 N -                   TCS IIF_QueryParameter "ecs.instrCool.tempChan310" skipped
2010/03/03 07:34:21.036867 N B                   telescope operations completed
2010/03/03 07:34:21.036923 W B                   recovered error "unknown" [src/lbc/actions.c:1603]
2010/03/03 07:34:21.036949 N B                   completed -45 OB image/s
2010/03/03 07:34:21.037135 N B                   completed -45 OB image/s
2010/03/03 07:34:21.037411 N B                   TCS IIF_QueryParameter "pcs.pointingStatus.target.parallacticAngle" skipped
2010/03/03 07:34:21.037456 N B                   telescope operations start
2010/03/03 07:34:21.037514 N B                   TCS waiting IIF_PresetTelescope or IIF_OffsetTelescope skipped
2010/03/03 07:34:21.037538 N -                   TCS IIF_QueryParameter "pcs.pointingStatus.achieved.achieved_RA.Radians" et al skipped
2010/03/03 07:34:21.037581 N B                   RA:-32768:-32768:-32768.-2147483648 DEC:+-32768:-32768:-32768.-2147483648
2010/03/03 07:34:21.037607 N B                   channel currently stopped
2010/03/03 07:34:21.140026 N -                   TCS IIF_QueryParameter "pcs.pointingStatus.achieved.achieved_RA.Radians" et al skipped
2010/03/03 07:34:21.140076 N B                   RA:-32768:-32768:-32768.-2147483648 DEC:+-32768:-32768:-32768.-2147483648
2010/03/03 07:34:21.250030 N -                   TCS IIF_QueryParameter "pcs.pointingStatus.achieved.achieved_RA.Radians" et al skipped
2010/03/03 07:34:21.250076 N B                   RA:-32768:-32768:-32768.-2147483648 DEC:+-32768:-32768:-32768.-2147483648
2010/03/03 07:34:21.360030 N -                   TCS IIF_QueryParameter "pcs.pointingStatus.achieved.achieved_RA.Radians" et al skipped
2010/03/03 07:34:21.360075 N B                   RA:-32768:-32768:-32768.-2147483648 DEC:+-32768:-32768:-32768.-2147483648

07:45 issued a couple LBC offsets from IRAF.

07:47 Play the OB again but LBC still in the funny state reporting "expd imgs -65/2"

07:51 lbckill.sh / lbcstart.sh

CA=-200

07:55 Play OB WT000350 x120 075549
This time guiding fails. From the old guider image, it looks like we may be too out of focus.
This time we uncheck the "use Blue" box on the LBCGUI, and that seems to stop the program from crashing.

dofpia, /redonly as LBTO on obs2

/home/lbcobs/LBCFPIA/lbcfpia//TCSSendWavefront/TCSSendWavefront: error while loading shared libraries: libTCSiif.so: cannot open shared object file: No such file or directory

Michele says this is because builds later than B35 have been linked statically rather than dynamically. This older LBC executable is still looking for the old library.

Increase Zglobal by 0.9

dofpia, /redonly as LBTO on obs2

Increase Zglobal by another 0.4

08:07 dofpia, /redonly as LBTO on obs2

08:12 Play OB WT000350 x120 081217 This is guiding OK.

Back to IRTC@LDG

08:20 pointto M5_1311 TelMode="ACQUIRE" gs=0 PAM- AzCam still having problems waiting for image to arrive. The movie image in ACQUIRE mode is intermittent. P.S. We still don't see any star.

08:33 play rfastextra OB (while IRTC is authorized on left) (but x120 was still set, so this is a goofy image) Move focus commands are succeeding on right side. But GetTrajectory for side "both" is rejected as expected.

Build BP1 tests with LBC

08:45 Michele switches us to Build BP1. John switches his TCSGUIs from obs2 to obs4. Some DD names have changes, so we'll ignore those errors.

09:05 Play wt000350 OB - We saw the star briefly at the correct location on IRTC, but it disappeared after a bit. Did Doug move the probe in front of it?

Loading a bootes OB crashes the LBC software

09:20 Play NDgrb1 (from grb080310) 092345 Guiding failed after a while because we are out of focus. Put my Zglobal offsets back in from earlier for the second exposure. 092759. Now we have 1.2 arcsec images (with no active optics) and we are guiding, but the guiding is not working very well. In the 4th image, it seems like the guiding is hopping back and forth between two stable positions.

09:44 Play NDgrb1 (from grb080310) - offset 0 2 in 094453 then offset 0 3 made the guiding fail. offset 0 2 then offset 2 0 in the second image 094847 second one made guiding fail.

Guider images are as good as 0.8 arcsec FWHM with no ACTIVE optics corrections of any kind except eyeball focus at midnight.

Back to IRTC@LDG with BP1

10:15 still couldn't find M5 star in ACQUIRE mode

10:25 Doug succeeds in finding Arcturus in IRTC in TRACK mode.

10:35 Set IE=-15 CA=48

10:37 pointto Arcturus TelMode=ACQUIRE gs=0 Never received any acquisition images

10:38 pointto Arcturus in TRACK mode

10:39 stopAGW -u 4, startAGW -u 4, home

10:48 pointto BS9155 Acquire No star

10:49 ptspiral Set IE=-3 CA=87

10:55 pointto BS9155 Active. Success. One Active Optics correction then Primary paniced and we lost our star

11:05 pointto BS9155 Active. Failed. Could not get full image from azcam. The error on the azcam server was that "could not send image size"

11:15 air pressure is low so Wayne will go check the air compressor

11:25 stop, startAGW, image acusition

11:34 Wayne fixes the popoff valve in the SX air box (using the standard technique of jiggling).

11:35 pointto BS9155 Active - Failed, no AGW selected. We had restarted GCS. setNewInstrument

11:37 pointto BS9155 Active - Success

11:42 Primary Mirror Panic

Some messing around with AzCam and presets here.

13:08 Close for impending dawn.

-- JohnHill - 03 Mar 2010
Topic revision: r5 - 04 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback